# Dynamic Scheduling for Vehicle-to-Vehicle Communications Enhanced Federated Learning

### v1

@@ VI-D Evaluation on Argoverse trajectory prediction dataset
Based on the dataset, the VFL system collaboratively trains a lane graph convolutional neural network (LaneGCN) [49]. The LaneGCN includes three sub neural networks: an ActorNet, a MapNet, and a FusionNet. The ActorNet contains a 1D CNN and a Feature Pyramid Network (FPN) to extract features of vehicle trajectories. The MapNet is a graph convolutional neural network that represents and extracts the map features. The FusionNet is used to fuse the vehicle trajectory features and the map features to output the final trajectory prediction results. We employ ADE as the metric for trajectory prediction, which is the average $l_{2}$ distance between the actual and predicted vehicle positions on the trajectory.
The performance of the proposed framework compared with the benchmarks is illustrated in Fig. 12. It is shown that the proposed VEDS algorithm outperforms the benchmarks both in terms of ADE. Specifically, VEDS achieves an ADE of $1.72$ after $2000$ rounds of training, which is $7.98\%$, $10.21\%$, $22.00\%$ lower than V2I-only, MADCA-FL and SA, respectively. These results validate the strong performance of our proposed VEDS algorithm when applied to real-world autonomous driving datasets.
@@ VII Conclusions
In this paper, we have considered a VFL system, where the SOVs and OPVs in a vehicular network collaborate to train an ML model under the orchestration of the RSU. A VEDS algorithm has been proposed to optimize the VFL training performance under energy constraints and channel uncertainty of vehicles. Convergence analysis has been performed to transform the implicit FL loss function into the number of successful aggregations. Then, a derivative-based drift-plus-penalty method has been proposed to convert the long-term stochastic optimization problem into an online MINLP problem, and a theoretical performance guarantee has been provided for the proposed transformation by bounding the performance gap between the online and offline solutions. Based on the analysis of the scheduling priority, the MINLP problem has been further reduced to a set of convex optimization problems, which can be efficiently solved using the interior-point method. Experimental results have illustrated that our proposed framework is robust under different vehicle speeds. The test accuracy is increased by $3.18\%$ for the CIFAR-10 dataset, and the ADE is reduced by $10.21\%$ for the Argoverse dataset.
@@ Appendix A Proof of Lemma 1
According to Assumption 3, there is
$\mathbb{E}\left[\frac{\sum_{m\in{\mathcal{S}}_{k}}I_{m,k}|\mathcal{D}_{m}|(\boldsymbol{w}_{m,k}-\boldsymbol{w}_{k-1})}{\sum_{m\in\mathcal{S}_{k}}I_{m,k}|\mathcal{D}_{m}|}\right]$
$=$ $\mathbb{E}\left[\frac{\sum_{m\in{\mathcal{S}}_{k}}\sum_{\boldsymbol{x}\in\mathcal{B}_{m,k}}I_{m,k}|\mathcal{D}_{m}|\eta_{k}\nabla f\left(\boldsymbol{w}_{k-1};\boldsymbol{x}\right)}{B_{k}\sum_{m\in\mathcal{S}_{k}}I_{m,k}|\mathcal{D}_{m}|}\right]$
$=$ $\eta_{k}\nabla F(\boldsymbol{w}_{k-1}).$
@@
${\nabla F(\boldsymbol{w}_{k-1}),\boldsymbol{w}_{k}-\boldsymbol{w}_{k-1}}\right>$ (31)
$=\big{<}{\nabla F(\boldsymbol{w}_{k-1}),\frac{\sum_{m\in{\mathcal{S}}_{k}}I_{m,k}|\mathcal{D}_{m}|(\boldsymbol{w}_{m,k}-\boldsymbol{w}_{k-1})}{\sum_{m\in\mathcal{S}_{k}}I_{m,k}|\mathcal{D}_{m}|}}\big{>}$
$=\eta_{k}\nabla\left\|{F(\boldsymbol{w}_{k-1})}\right\|^{2}.$
@@
$\leq\left\|{\nabla F(\boldsymbol{w}_{k-1})}\right\|^{2}+\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}I_{m,k}}.$ (32)
@@ Appendix B Proof of Theorem 1
According to Assumption 2 ($\mu$-strong convexity of the loss functions), the Polyak-Lojasiewicz inequality holds
@@
$\mathbb{E}[F({\boldsymbol{w}}_{K})]-F(\boldsymbol{w}^{*})$
$\leq(1-\mu\eta_{K})(\mathbb{E}[F({\boldsymbol{w}}_{K-1})]-F(\boldsymbol{w}^{*}))+\frac{\eta_{K}}{2}\frac{G^{2}}{B_{K}\sum_{m\in\mathcal{S}_{K}}I_{m,K}}$
@@ Appendix C Proof of Theorem 2
Based on the definition of $q_{m}^{\text{SOV}}(t)$, we have $q_{m}^{\text{SOV}}(t+1)-q_{m}^{\text{SOV}}(t)\leq\phi_{m}^{\text{SOV}}(t),\forall m\in\mathcal{S}_{k},t\in\mathcal{T}_{k}$, and therefore
@@
Appendix D Proof of Proposition 1
@@
$\mathcal{L}=V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\kappa\beta\log_{2}\left(1+\frac{p_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}\right)$
$-\kappa q_{m}^{\text{SOV}}(t)p_{m}(t)-\lambda_{m}p_{m}(t)+\nu_{m}(p^{\text{max}}_{m}-p_{m}(t)).$
@@
$\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\kappa\beta\frac{|h_{m,r}(t)|^{2}}{\beta N_{0}}}{1+\frac{p_{m}^{*}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}}-q_{m}^{\text{SOV}}(t)\kappa-\lambda_{m}^{*}+\nu_{m}^{*}=0.$
@@
$p_{m}^{*}(t)=\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\beta}{q_{m}^{\text{SOV}}(t)}-\frac{\beta N_{0}}{|h_{m,r}(t)|^{2}}.$
@@
$p_{m}^{*}(t)=\left[\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\beta}{q_{m}^{\text{SOV}}(t)}-\frac{\beta N_{0}}{|h_{m,r}(t)|^{2}}\right]^{p^{\text{max}}_{m}}_{0},$

### v2

@@ VI-B5 Impact of Model Size
Figure 14: Test accuracy on the CIFAR-10 dataset compared with frequency selective channel and highway scenarios (non-i.i.d. setting).
@@
VI-C Evaluation on the CIFAR-10 Dataset
Then, we evaluate the proposed VEDS algorithm on the CIFAR-10 dataset[56], which comprises $50000$ training images and $10000$ test images across ten categories. We consider both the independent and identically distributed (i.i.d.) and non-independent and identically distributed (non-i.i.d) settings. For the i.i.d. setting, the dataset is evenly divided into $40$ subsets, each containing samples from all 10 categories. For the non-i.i.d. setting, data samples are organized by category, and each vehicle holds a disjoint subset of data with samples from $2$ categories. Using the dataset, we train a convolutional neural network (CNN). This CNN includes six convolutional layers. Three of them are followed by a ReLU activation layer and a max pooling layer, and the other three are followed by a ReLU activation layer and a normalization layer. The last convolutional layer is connected to a fully connected layer and a softmax output layer for classification. The batch size of each vehicle is randomly chosen from 16, 32 and 48, and the learning rate is set to 0.1.
The test accuracy of the VEDS algorithm compared with the benchmarks is illustrated in Fig. 15 (i.i.d.) and Fig. 15 (non-i.i.d.), where $\alpha=2,V=0.2$, and the maximum vehicle speed $v^{\text{max}}=10$ m/s. In the i.i.d. scenario, both VEDS and the benchmarks achieve high test accuracy. The VFL convergence speed of the VEDS algorithm closely approaches that of the optimal benchmark and is significantly higher than other benchmarks. Under the non-i.i.d. scenario, the convergence speed and the highest test accuracy of the VEDS algorithm are close to the optimal benchmark and exceed the other three benchmarks. After $1000$ rounds of training, VEDS achieves a test accuracy of $62.42\%$, exceeding V2I-only, VEFL, MADCA-FL and SA over $10.57\%$, $19.41\%$ $26.90\%$ and $41.42\%$. After $10000$ rounds of training, the highest achievable accuracies are $79.84\%$ for the optimal benchmark, $78.78\%$ for VEDS, $75.80\%$ for V2I-only, $75.61\%$ for VEFL, $74.74\%$, and $73.23\%$ for SA.
We also evaluate the performance of the VEDS algorithm under frequency selective channels (denoted as VEDS-F), as discussed in Section V-G. Additionally, we extend the evaluation to highway scenarios, where the pathloss models for LOS and NLOSv channels are modified by $PL_{\text{LOS}}=38.77+16.7\log_{10}{d}+18.2\log_{10}{\gamma}$. Fig. 15 shows that the highest testing accuracy in highway environments is slightly better, exceeding that in urban environments by $0.16\%$. For the frequency selective channel scenario, we use the average channel condition to estimate the actual channel state, which introduces a slight performance degradation. While the convergence speed of the VEDS algorithm under frequency selective fading channels is slightly lower than that under frequency non-selective fading channels, the highest testing accuracy is well-preserved. The highest testing accuracy under the frequency selective fading scenario reaches $78.44\%$ for the highway case and $78.28\%$ for the urban case, which is $0.60\%$ and $0.64\%$ lower than the frequency non-selective fading scenario, respectively. These results demonstrate the applicability of the VEDS algorithm under various vehicular environments and channel conditions.
@@ VI-D Evaluation on Argoverse Trajectory Prediction Dataset
Based on the dataset, the VFL system collaboratively trains a lane graph convolutional neural network (LaneGCN) [61]. The LaneGCN includes three sub-neural networks: an ActorNet, a MapNet, and a FusionNet. The ActorNet contains a 1D CNN and a Feature Pyramid Network (FPN) to extract features of vehicle trajectories. The MapNet is a graph convolutional neural network that represents and extracts the map features. The FusionNet is used to fuse the vehicle trajectory features and the map features to output the final trajectory prediction results. The batch size of each vehicle is randomly chosen from 16, 32 and 48, and the learning rate is set to 0.1. We employ ADE as the metric for trajectory prediction, which is the average $l_{2}$ distance between the actual and predicted vehicle positions on the trajectory.
The performance of the proposed framework compared with the benchmarks is illustrated in Fig. 15. It is shown that the proposed VEDS algorithm outperforms the benchmarks both in terms of ADE. Specifically, VEDS achieves an ADE of $1.72$ after $2000$ rounds of training, which is $7.14\%$, $9.82\%$, $16.30\%$ and $22.09\%$ lower than V2I-only, VEFL, MADCA-FL and SA, respectively. These results validate the strong performance of our proposed VEDS algorithm when applied to real-world autonomous driving datasets.
@@ VII Conclusions
This paper has considered a VFL system, where the SOVs and OPVs in a vehicular network collaborate to train an ML model under the orchestration of the RSU. A VEDS algorithm has been proposed to optimize the VFL training performance under energy constraints and channel uncertainty of vehicles. A convergence analysis has been performed to transform the implicit FL loss function into the number of successful aggregations. Then, a derivative-based drift-plus-penalty method has been proposed to convert the long-term stochastic optimization problem into an online MINLP problem, and a theoretical performance guarantee has been provided to bound the performance gap between the online and offline solutions. Based on the analysis of the scheduling priority, the MINLP problem has been further reduced to a set of convex optimization problems, solved by the interior-point method. Experimental results have demonstrated the robustness of the framework under varying vehicle speeds. The test accuracy is increased by $4.20\%$ for the CIFAR-10 dataset, and the ADE is reduced by $9.82\%$ for the Argoverse dataset.
In the future, we plan to extend our current framework by incorporating sparsification and quantization to reduce the communication overhead. Furthermore, we aim to develop scheduling strategies that consider the diverse distributions of local datasets across different to increase the model robustness and generalization. We will also explore the application of our proposed method on a broader range of datasets, including tasks such as autonomous driving object detection and semantic segmentation. These future developments will address current limitations and expand the practical applicability of our framework to complex and diverse scenarios.
@@ Appendix A Proof of Lemma 1
$\mathbb{E}\left[\left<{\nabla F(\boldsymbol{w}_{k-1}),\boldsymbol{w}_{k}-\boldsymbol{w}_{k-1}}\right>\right]$
$=\mathbb{E}\left[\big{<}{\nabla F(\boldsymbol{w}_{k-1}),\frac{\sum_{m\in{\mathcal{S}}_{k}}I_{m,k}(\boldsymbol{w}_{m,k}-\boldsymbol{w}_{k-1})}{\sum_{m\in\mathcal{S}_{k}}I_{m,k}}}\big{>}\right]$
$=-\mathbb{E}\left[\big{<}{\nabla F(\boldsymbol{w}_{k-1}),\frac{\sum_{m\in{\mathcal{S}}_{k}}I_{m,k}\eta_{k}\nabla f\left(\boldsymbol{w}_{k-1},\mathcal{B}_{m,k}\right)}{\sum_{m\in\mathcal{S}_{k}}I_{m,k}}}\big{>}\right]$
$=-\eta_{k}\left\|{\nabla F(\boldsymbol{w}_{k-1})}\right\|^{2}.$ (32)
@@
$\leq\eta_{k}^{2}\left\|{\nabla F(\boldsymbol{w}_{k-1})}\right\|^{2}+\frac{\eta_{k}^{2}G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}I_{m,k}}.$ (33)
@@ Appendix B Proof of Theorem 1
Assuming that the loss function is the loss $\mu$-strongly convexity, the Polyak-Lojasiewicz inequality holds
@@
$\mathbb{E}[F({\boldsymbol{w}}_{K})]-F(\boldsymbol{w}^{*})\leq(1-\mu\eta_{K})(\mathbb{E}[F({\boldsymbol{w}}_{K-1})]-F(\boldsymbol{w}^{*}))$
$+\frac{\eta_{K}}{2}\frac{G^{2}}{B_{K}\sum_{m\in\mathcal{S}_{K}}I_{m,K}}$
@@ Appendix C Proof of Theorem 2
Based on lemma 1, and setting $\eta_{k}\leq\frac{1}{L}$, we have
$\frac{\eta_{k}}{2}\mathbb{E}\left\|\nabla F(\boldsymbol{w}_{k-1})\right\|^{2}$
$\leq\mathbb{E}[F(\boldsymbol{w}_{k-1})]-\mathbb{E}[F(\boldsymbol{w}_{k})]+\frac{L\eta_{k}^{2}}{2}\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}I_{m,k}}.$
Taking a telescopic sum from $k=1$ to $k=K$, and setting $\eta_{k}=\frac{1}{K^{1/2}L}$ we get
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left\|\nabla F(\boldsymbol{w}_{k-1})\right\|^{2}$
$\leq\frac{2L\left(F(\boldsymbol{w}_{0})-\mathbb{E}[F(\boldsymbol{w}_{K})]\right)}{K^{1/2}}+\frac{1}{K^{3/2}}\sum_{k=1}^{K}\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}I_{m,k}}$
$\leq\frac{2L\left(F(\boldsymbol{w}_{0})-F(\boldsymbol{w}^{*})\right)}{K^{1/2}}+\frac{1}{K^{3/2}}\sum_{k=1}^{K}\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}I_{m,k}}.$
This completes the proof.
Appendix D Proof of Theorem 3
@@
Based on the definition of $q_{m}^{\text{SOV}}(t)$, we have $q_{m}^{\text{SOV}}(t+1)-q_{m}^{\text{SOV}}(t)\leq\phi_{m}^{\text{SOV}},\forall m\in\mathcal{S}_{k},t\in\mathcal{T}_{k}$, and therefore
@@
Appendix E Proof of Proposition 1
@@
$\mathcal{L}=V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\kappa\beta\log_{2}\left(1+\frac{p_{m}(t)h_{m,r}(t)}{\beta N_{0}}\right)$
$-\kappa q_{m}^{\text{SOV}}(t)p_{m}(t)+\lambda_{m}p_{m}(t)+\nu_{m}(p^{\text{max}}_{m}-p_{m}(t)).$
@@
$\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\kappa\beta\frac{h_{m,r}(t)}{\beta N_{0}}}{1+\frac{p_{m}^{*}(t)h_{m,r}(t)}{\beta N_{0}}}-q_{m}^{\text{SOV}}(t)\kappa+\lambda_{m}^{*}-\nu_{m}^{*}=0.$
@@
$p_{m}^{*}(t)=\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\beta}{q_{m}^{\text{SOV}}(t)}-\frac{\beta N_{0}}{h_{m,r}(t)}.$
@@
$p_{m}^{*}(t)=\left[\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\beta}{q_{m}^{\text{SOV}}(t)}-\frac{\beta N_{0}}{h_{m,r}(t)}\right]^{p^{\text{max}}_{m}}_{0},$